Minutes of the code management group meeting 11/12/2014

From UKCA
Jump to: navigation, search

Agenda

  1. Minutes of the last meeting
  2. UKCA Improvements for UKESM1
  3. vn9.2/GA7.0 update
  4. Release jobs
  5. Rose stem jobs
  6. Bug fixing
  7. Targets for next release (10.1)
  8. Shared repository
  9. A.O.B.

Present: Zak Kipling (ZK), Luke Abraham (LA), Nicolas Bellouin (NB), Graham Mann (GM), Sandip Dhomse (SD), Mark Richardson (MR), Fiona O'Connor (FOC), Alan Hewitt (AH), Colin Johnson (CJ), Nick Savage (NS), Jo Anderson (JA), Jane Mulcahy (JM), Mohit Dalvi (MD)

Minutes of the last meeting

  • (1) NS to circulate to all those with access to the mailbox information on how to access ukca_bugs emails through Outlook. Done - see also item 6
  • (2) ALL. Communicate any foreseen UKCA improvements to FOC before next Code Mgmt meeting. Done
  • (3) JM to check whether settings, branches in vn8.4 Release candidate are consistent with GA6/7 assessment job and also attempt tuning of SO2 scavenging parameters. In progress - liase with LA?
  • (4) NS/LA. Add standing agenda item on Bug Reporting. Done
  • (5) NS to email Nicolas B. regarding participation in CMG meetings/ process. Done
  • (6) NS. Create doodle poll for next meeting. Done

Actions Carried forward from Last Meeting

  • (1) CJ to circulate plots of radical concentrations in stratosphere showing odd banded structure, including version with different photolysis scheme options. Ongoing. Done a while ago - nothing much has changed?
    • Action: FOC to put on UKESM1 document
  • (2) CJ to email LA re problems with SO2 diagnostics for emissions. Issues related to using ASAD diagnostics package. Ongoing - needs debugging.
    • Action: CJ to put branch on PUMA email CMG location of branch to check details
  • (3) CJ to open ticket and start work on resolving inconsistencies in aerosol chemistry schemes. (unused tracers-CS2, missing chemistry in CHEST, above diagnostics) Ongoing. Done vn8.6 - see pt.1 of UKESM1 document
  • (4) GM check 8.4 versus 7.3 GLOMAP-mode code and report any bug fixes to Mohit. Ongoing. Still Ongoing (big task) - not just bugfixes, also functionality updates

UKCA Improvements for UKESM1

Discussion centred around document circulated by FOC.

  • Action: FOC to add in status and ticket numbers to each item
  • Action: FOC to reorder by priority

Discussion on individual items:

  • (4) Dry deposition: split into 2 options:
    1. NS "quick and dirty fix" first (FOC)
    2. AH for future proof.
Don't anticipate changes to gas-phase species.
  • (5) Not sure about priority. Could create large changes - may not be able to trial to get in for GA7.
  • (6) Risk L
  • (8) Risk L as can use prescribed
  • (9)
    • Action: FOC/NB/GM - discuss (9) offline
  • Action: FOC to talk to Colin Jones about nitrate (high priority), with a offer to help Steve Rumbold
  • Action: JM to make standing telecon from January to discuss this list. Discuss with Alistair Sellar
  • Action: ALL - feed in to FOC list of priorities

vn9.2/GA7.0 update

Package testing for GA7. Some issues. Aerosol loading high over anthropgenic regions. There is a negative impact in terms of biases. JM/CJ looking at SO2 to tune this down. When implemented with cloud/rad changes this leads to a large increase in cloud liquid water which then further degrades radiation biases. This is currently being investigated. The needs to be some Cloud/Aerosol tuning, and this is not straight-forward as there are complicated interactions with cloud.

  • Action: JM to report back on progress at the next meeting

The vn8.4 RC6.2 has this issue as well. Surface values are in reasonable agreement (SO2 high over Europe, but no effect on aerosol).

  • Action: JM/GM to talk offline

The wet deposition is tunable.

Release jobs

See also Vn8.4 GA4.0 Release Candidate: RC6.2.

The aerosol column loading too high, but concentrations. at surface are in good agreement. This is a Persistent problem and needs tuning - related to GA7.0 (above)

Aerosol

The GLOMAP-mode output is scientifically usable. Can be released as is now. Surface properties green. Column properties amber.

Chemistry

Chemistry - happy with caveats. Amber for both troposphere and stratosphere.

  • Action: LA to release RC6.2

Rose stem jobs

Updates for the UKCA Rose stem jobs are in plan for 10.x. Needs branch and change to rose meta. Some issues with age of air job.

The available apps will be:

  • proto GA7 (nightly)
  • proto UKESM1 (weekly)

Rose stem needs to be forward looking, but needs to use lodged code

  • Action: MD to do for 10.2 testing (needs updates to 10.1)

Note: rose stem is site specific and may only be available on MONSooN. If this isn't the case we may need to contact the JWCRP.

  • Action: LA to talk to Ros Hatcher at NCAS-CMS about NCAS plans for Rose stem.

Bug Reporting

See also Bugfixes (new approach).

  • Action: Incorporate Bug #35 into dry deposition work (see above).
  • Action: Bug #36 (ozone): FOC/Steven Hardiman to take on

There are two elements to this list:

  1. Monitoring the bug list.
  2. Having owners to resolve each bug.

It was decided to try to have the responsibility for fixing bugs circulate between different institutes, to try to avoid too much falling to one person/group.

  • Action: CMG to monitor list of bugs and circulate between group for owners
  • Action: At CMG future meetings there should be included an appendix listing the open bugs. If a ticket is included it is not an open bug (as the owner of ticket is person doing lodging).

Targets for next release (10.1)

  • Next UM code freeze: 13th February 2015
  • 10.1 released: 27th February 2015 (or earlier)
  • High - Offline oxidants BE solver, SO2 diags, H2O treatment
  • AQUM - Carlos Ordonez is looking at ensuring iBVOC works with netCDF emissions
  • FOC - bugfix for O3 (Bug #36 - see above)
  • MD - rose stem required updates

Shared repository

See also Met Office Science Repository Service (MOSRS).

There is no automatic migrating of tickets over from the internal Met Office trac pages.

  • Action: ALL to migrate tickets over

A.O.B.

  • Action: LA to set up doodle poll for next meeting for after the code freeze - late Feb/Early March.